# Random forest–based out-of-distribution detection for robust lung cancer auto-segmentation

### v2

@@ Abstract
Accurate detection and segmentation of cancerous lesions from computed tomography (CT) scans is essential for automated treatment planning and cancer treatment response assessment. Transformer-based models with self-supervised pretraining have achieved strong performance on in-distribution (ID) data but often generalize poorly on out-of-distribution (OOD) inputs. We investigate this behavior for lung cancer segmentation using an encoder–decoder model. Our encoder is a Swin Transformer pretrained with masked image modeling (SimMIM) on 10,432 unlabeled 3D CT scans spanning cancerous and non-cancerous conditions, and the decoder was randomly initialized. This model was evaluated on an independent ID test set and four OOD scenarios: chest CT cohorts (pulmonary embolism and negative COVID-19) and abdomen CT cohorts (kidney cancers and non-cancerous pancreas). OOD detection was performed at the scan level using RF-Deep, a random forest classifier applied to contextual tumor-anchored feature representations. We evaluated 920 3D CTs (172,650 images) and observed that RF-Deep achieved FPR95s of 18.26% and 27.66% on the chest CT cohorts, and near-perfect detection (<< 0.1% FPR95) on abdomen-CT cohorts, consistently outperforming established OOD methods. These results demonstrate that our RF-Deep classifier provides a simple, lightweight and effective approach for enhancing the reliability of segmentation models in clinical deployment.
@@
Swin Transformer, auto-segmentation, lung cancer, out-of-distribution
@@ 1 Purpose
Among deep learning architectures, models that combine pretrained transformer backbones with convolutional decoders have demonstrated high accuracies in volumetric auto-segmentation of organs and tumors from 3D radiological scans [Willemink2022, jiang2022self, NguyenAAAI2023, YanWACV2023, Qayyum2023BHI, gu2024build]. A major challenge in deploying these models at scale is the potential accuracy degradation when applied to real-world scenarios that differ from those seen during training. For example, models trained to segment malignant lung cancers from chest CTs may also encounter scans that have benign nodules from lung cancer screening or scans containing unrelated diseases such as pulmonary embolisms, producing unanticipated and incorrect results outside their intended scope.
@@
Traditional segmentation metrics such as the Dice similarity coefficient (DSC) and Hausdorff distance at 95th percentile (HD95) are designed to assess performance on in-distribution (ID) datasets, but are insufficient indicators of models’ robustness to out-of-distribution (OOD) data. OOD detection methods relying on model confidence scores, such as MaxSoftmax [hendrycks17baseline], often fail in scenarios where models produce confidently incorrect segmentations [yeung2023calibrating, hendrycks2019anomalyseg, zimmerer2022mood]. Alternative approaches using secondary models like VQ-GANs [graham2022transformer, pinaya2022unsupervised] require large training data, are computationally intensive, and lack interpretability, limiting their practical utility in high-throughput clinical workflows. Radiomics feature-based methods [vasiliuk2023limitations, konz2024radmetricmedicalimage] have shown to be effective at distinguishing scans of organs but have not generalized yet to tumors. To address these limitations, we propose RF-Deep, a lightweight scan-level OOD detection framework that leverages tumor-anchored contextual features from backbone encoders for reliable clinical deployment.
We evaluated our approach on far-OOD scans (diseases in different anatomic sites) and challenging near-OOD scans (different diseases in the same site with similar appearance). Our contributions are:
  1. i.
    a lightweight, model–agnostic OOD detection framework using random forest on deep feature representations extracted from the segmentation model’s encoder. While random forests for OOD detection are established, our contribution lies in the task-anchored application, a practical deployment concern under-explored in medical imaging,
  2. ii.
    demonstration on lung cancer segmentation including near-OOD (pulmonary embolism, COVID-19) and far-OOD (kidney cancers, healthy pancreas) scenarios, and
  3. iii.
    a systematic comparison against established OOD detection methods using five public datasets comprising 172,650 images from 920 patient scans.
@@ 2.1 OOD detection task definition
Let $\mathcal{D}_{\text{in}}$ denote the distribution of lung cancer CT scans, closely matching the lung CT data used to fine-tune our segmentation model, and $\mathcal{D}_{\text{out}}$ represents the scans that differ in pathology and anatomy. Our objective is to determine whether a new scan $x$ belongs to $\mathcal{D}_{\text{in}}$ or $\mathcal{D}_{\text{out}}$, at the scan level. We compute voxel-level metric scores, average them to obtain $S(x)$, and apply a decision threshold $\tau$: scans with $S\leq\tau$ are flagged as ID, and $S\geq\tau$ as OOD. Since tumor segmentation is binary, aggregation is restricted to the predicted tumor regions, leveraging model-relevant spatial context while excluding irrelevant anatomy.
2.2 Segmentation model architecture
@@
Our segmentation model is a hybrid transformer–convolution architecture that combines the global contextual modeling of a hierarchical Swin Transformer [liu2021swin] encoder with the local spatial precision of a U-Net–based [ronneberger2015u] convolutional decoder to accurately delineate anatomical boundaries of tumors. The encoder uses a depth configuration of 221222-2-12-2 across four stages with $4\times 4\times 4$ patch size, enabling multi-scale feature aggregation, while windowed self-attention reduces computational complexity yet preserves global context for large volumetric inputs (128 $\times$ 128 $\times$ 128 voxels). The decoder was randomly initialized, whereas the encoder was pretrained using SimMIM [xie2022simmim], a self-supervised masked image modeling approach in which 75% of 3D patches are masked and reconstructed. This strategy leverages large-scale unlabeled CT data to learn robust, anatomy-aware feature representations [hatamizadehSwinUNETR2022, tang2022self, jiang2022self, jiang2024self] and provides a strong foundation for downstream tumor segmentation and OOD detection.
2.3 OOD Detection with RF-Deep
Patient scans exhibit high variability and unlike natural images, may span the large anatomical boundaries, making a global OOD assessment inefficient. We therefore focus on the most semantically relevant areas, the tumor regions predicted by the segmentation model. Our approach (Fig. 1) can be outlined in four steps: (i) fine-tune a segmentation model with pretrained backbone encoders for lung tumor segmentation, (ii) extract tumor-anchored 3D crops and multi-scale encoder features, (iii) aggregate features into fixed-length representation maintaining global context focused on clinically relevant tumor regions and train a random forest classifier (RF-Deep) with ID and OOD examples, and (iv) classify new scans as ID or OOD using the trained RF-Deep. Critically, RF-Deep adds negligible computational cost as the feature extraction stage leverages the already-stored encoder representations from segmentation, and random forest inference is orders of magnitude faster than running secondary deep models.
2.4 Implementation details
@@
Datasets. We fine-tuned the model on NSCLC Radiomics [aerts2015data] (N=317), with LRAD [bakr2017data] (N=140) as the ID dataset. OOD evaluation used RSNA-Pulmonary Embolism [colak2021rsna] (N=120), negative COVID-19 MIDRC [ricord2021covid] (N=120), KiTS-23 [heller2023kits21] (N=120), and Pancreas-CT [roth2015deeporgan] (N=82), totaling 920 scans.Segmentation. All models were implemented using PyTorch [paszke2019pytorch] and MONAI [cardoso2022monai], fine-tuned using cross-entropy and Dice loss with a batch size of 16 across 4 NVIDIA GPUs. We used a learning rate of $2\times 10^{-4}$, with linear warm-up and cosine annealing over 1000 epochs. Augmentations included flips, rotations, affine, and intensity shifts. Inputs were normalized (HU [400,400][-400,400]), resampled (1mm3), and cropped to $128^{3}$. Sliding window with 50% overlap was used for inference. RF-Deep. Random forest (1000 trees, max depth 20, balanced weights) via the scikit-learn library. Features were extracted from 8 tumor-anchored crops per scan, and predictions were averaged at inference. Comparative approaches. RF-Radiomics used 293 IBSI-compliant features [zwanenburg2020image] via PyCERR [deasy2003cerr, apte2018extension], reduced with recursive elimination, and with same config as RF-Deep. We also evaluated MaxSoftmax [hendrycks17baseline], MaxLogits [hendrycks2022logits], energy [liu2020energy], and entropy. OOD experiment protocol. RF-based approaches used a fixed 40/60 patient-level train–test split, repeated over 100 matched seeds to assess stability. Other baselines used the full cohorts as they require no auxiliary data. Evaluation metrics. We report AUROC and FPR95 (%) to measure ID–OOD separability [davis2006relationship, liang2017enhancing].
@@ 3 Results
Table 1: Out-of-distribution (OOD) detection performance across different OOD detection strategies for our segmentation model. Results are reported as AUROC ($\uparrow$) and FPR95 ($\downarrow$), with best values per dataset highlighted in bold.
Method Pulmonary embolism COVID-19 Kidney cancer Pancreas
AUROC FPR95 (%) AUROC FPR95 (%) AUROC FPR95 (%) AUROC FPR95 (%)
@@
Our results demonstrate that RF-Deep provides robust scan-level OOD detection for lung cancer segmentation, outperforming established approaches including MaxSoftmax and MaxLogits ( 1). Encoder features from the SimMIM-pretrained and fine-tuned model, paired with RF, consistently yielded higher AUROC and lower FPR95 across all scenarios, with large margin gains in the abdominal cohorts (FPR95 $\leq$ 0.1, second best is 6 points away). The substantial performance differences, particularly in far-OOD scenarios, reflect the advantage of task-relevant contextual features over confidence-based heuristics. Fig. 2 visually shows RF-Deep’s ability to detect scans from completely different anatomical sites and different disease cases, with some limitations. In pulmonary embolism cases, we observed that the segmentation model often highlights tumor-like structures rather than the emboli; nevertheless, our contextual features enable RF-Deep to correctly classify these scans as OOD in most cases, surpassing non-contextual approaches. Additionally, we computed an unsupervised clustering of the features [masarczyk2024tunnel] across all cohorts (Fig. 3(a)) and observed semi-clear, distinct clusters for in-distribution (ID) and OOD cohorts, making them separable with the random forest classifier. Fig. 3(b) and Fig. 3(c) show the SHAP analysis on the RF-Deep providing interpretability and identifying which encoder features are most influential in classifying a scan as OOD, allowing for future work into feature reduction. Finally, stage-wise ablation (Fig. 3(d)), wherein only features corresponding to a particular stage of the transformer were used for training RF-Deep, demonstrated that mid- and early-deep level Swin Transformer features are most effective, aligning with our SHAP findings.

### v3

# Random forest-based out-of-distribution detection for robust lung cancer segmentation

@@ Abstract
Accurate detection and segmentation of cancerous lesions from computed tomography (CT) scans is essential for automated treatment planning and cancer treatment response assessment. Transformer-based models with self-supervised pretraining can produce reliably accurate segmentation from in-distribution (ID) data but degrade when applied to out-of-distribution (OOD) datasets. We address this challenge with RF-Deep, a random forest classifier that utilizes deep features from a pretrained transformer encoder of the segmentation model to detect OOD scans and enhance segmentation reliability. The segmentation model comprises a Swin Transformer encoder, pretrained with masked image modeling (SimMIM) on 10,432 unlabeled 3D CT scans covering cancerous and non-cancerous conditions, with a convolution decoder, trained to segment lung cancers in 317 3D scans. Independent testing was performed on 603 3D CT public datasets that included one ID dataset and four OOD datasets comprising chest CTs with pulmonary embolism (PE) and COVID-19, and abdominal CTs with kidney cancers and healthy volunteers. RF-Deep detected OOD cases with a FPR95 of 18.26%, 27.66%, and << 0.1% on PE, COVID-19, and abdominal CTs, consistently outperforming established OOD approaches. The RF-Deep classifier provides a simple and effective approach to enhance reliability of cancer segmentation in ID and OOD scenarios.
@@
Swin, lung cancer segmentation, out-of-distribution detection
@@ 1 Purpose
Pretrained transformer based deep learning (DL) methods combined with convolutional decoders have demonstrated capability to segment organs and tumors from radiological images [Willemink2022, jiang2022self, NguyenAAAI2023, YanWACV2023, Qayyum2023BHI, gu2024build]. A major challenge in deploying DL segmentation models at scale in research and clinical settings is the potential accuracy degradation when the same models are applied to real world scenarios that differ from those seen during training. For example, models trained to segment malignant lung cancers from chest CTs for treatment planning may also be applied to scans with benign nodules from lung cancer screening or scans containing unrelated diseases such as pulmonary embolisms. Although all of these involve chest CTs, models trained for cancer segmentation can produce unanticipated and incorrect results when applied outside their intended scope.
@@
Traditional segmentation evaluation metrics such as the Dice similarity coefficient (DSC) and Hausdorff distance at 95th percentile (HD95) are designed to assess model performance on in-distribution (ID) datasets, but are insufficient indicators of the model’s robustness to out-of-distribution (OOD) data. OOD detection approaches relying on model confidence scores, such as MaxSoftmax [hendrycks17baseline], often fail in scenarios where models produce confidently incorrect segmentations, occurring in both medical [yeung2023calibrating] and natural image analyses [hendrycks2019anomalyseg]. These methods have been applied to distinctly different disease sites such as the lung and abdomen [mehrtash2020confidence, zimmerer2022mood], a relatively easier task compared to detecting OOD cases within the same site. Alternative approaches using secondary models, such as VQ-GANs [graham2022transformer, pinaya2022unsupervised], require substantially large secondary training data, are computationally intensive, and lack interpretability, limiting their practical utility in high-throughput clinical workflows. Finally, radiomics feature-based methods that leverage standardized features have been shown to be effective at distinguishing scans of organs but have not generalized to tumors [vasiliuk2023limitations, konz2024radmetricmedicalimage]. To address these limitations, we developed a random forest classifier combining deep features (RF-Deep) of a model trained to segment lung cancers to detect OOD scans.
We evaluated our OOD detection approach on both, far-OOD scans consisting of diseases occurring in different anatomic site, and the more challenging near-OOD scans where different diseases occur in the same anatomic site with similar appearance statistics as the training datasets. In summary, our contributions are: (a) a lightweight, random forest classifier based OOD detection framework utilizing deep features trained to extract useful feature representations for ID cases. The deep features use the encoder features in order to leverage the robustness of the pretrained encoder to imaging distribution variations inherent even in ID datasets. The classifier is trained on a small set of ID and representative OOD examples, following the outlier exposure paradigm [hendrycks2018deep], (b) a demonstration of the effectiveness of our approach on lung cancer segmentation from 3D CT scans, including both near-OOD scenarios in chest CTs (pulmonary embolism and COVID-19) and far-OOD scenarios in abdominal CTs (kidney cancers and non-cancerous pancreas), and (c) a systematic comparison of our RF-Deep approach against common OOD detection methods (like MaxSoftmax) and a radiomics feature–based classifier (RF-Radiomics), using five public datasets of 603 patient scans.
@@ 2 Method
OOD detection task definition. Let $\mathcal{D}_{\text{in}}$ denote the distribution of lung cancer CT scans, closely matching the training dataset used to create the lung tumor segmentation model, and $\mathcal{D}_{\text{out}}$ represents the scans that differ in pathology and anatomic site. The objective is to determine whether a new scan $x$ belongs to $\mathcal{D}_{\text{in}}$ or $\mathcal{D}_{\text{out}}$, at the scan level, via a scoring mechanism $S(x)$. Since tumor segmentation is binary, aggregation is restricted to the predicted tumor regions, leveraging model-relevant spatial context while excluding irrelevant anatomy.
@@
Segmentation model architecture. A hybrid transformer encoder-convolutional decoder architecture was employed. The encoder extracts global and spatially local contexts through a hierarchical Swin Transformer [liu2021swin] encoder. The decoder leverages local spatial precision of a U-Net–based [ronneberger2015u] convolutional network to accurately delineate anatomical boundaries. The encoder uses a depth configuration of 221222-2-12-2 across four stages with $4\times 4\times 4$ patch size, enabling multi-scale feature aggregation, while windowed self-attention reduces computational complexity yet preserves global context for large volumetric inputs (128 $\times$ 128 $\times$ 128 voxels). Encoder was initialized with pretrained weights with self-supervised learning performed using SimMIM [xie2022simmim] method that uses masked image modeling (MIM) approach to predict and reconstruct masked image patches. Pretraining used 10,412 unlabeled 3D CT scans sourced from public and institutional datasets [jiang2022self, jiang2024self] with MIM task performed by randomly masking 75% of 3D patches in the image. Note that SSL pretraining is a unsupervised pretraining approach that does not require labeled image datasets for pretraining. Pretrained encoder was used to extract features that are robust to CT image acquisition features to accurately identify ID cases despite imaging differences and detect OOD cases.
OOD Detection with RF-Deep classifier. Unlike object centric photographic images, medical images may contain varying anatomic extent, making a global image-based OOD assessment inefficient and potentially less accurate. Hence, we leveraged the target task-relevant regions or tumors as extracted by the segmentation model to focus OOD detection on relevant image regions. Our approach (Fig. 1) consists of four steps: (i) fine-tune a segmentation model for lung tumor segmentation, (ii) obtain tumor-centered 3D image regions, (iii) extract an aggregate feature representation within the generated tumor regions from the multi-scale transformer encoder layers and train a RF-Deep classifier using the deep features using ID and OOD examples, and (iv) the extracted RF-Deep classifier can then be used to distinguish ID from OOD samples.
@@
Implementation details. The segmentation model was trained from a public dataset containing non-small cell lung cancers [aerts2015data] (N=317). It was implemented using PyTorch [paszke2019pytorch] and MONAI [cardoso2022monai], fine-tuned using cross-entropy and Dice loss with a batch size of 16 across 4 NVIDIA GPUs. A learning rate of $2\times 10^{-4}$ was used, with linear warm-up and cosine annealing over 1000 epochs. Augmentations included flips, rotations, affine, and intensity shifts. Inputs were normalized (HU [400,400][-400,400]), resampled (1mm3), and cropped to $128^{3}$. Sliding window with 50% overlap was used for inference. RF-Deep classifier consisted of a random forest (1000 trees, max depth 20, balanced weights) trained via the scikit-learn library. Features were extracted from 8 tumor-centered crops per scan, and predictions were averaged at inference. Experiments to detect OOD utilized a held-out ID dataset containing 140 lung cancers was used [bakr2017data] and four datasets consisting of pulmonary embolism (PE) [colak2021rsna] (N=120), COVID-19 [ricord2021covid] (N=120), kidney cancers [heller2023kits21] (N=120), and healthy abdominal CT scans [roth2015deeporgan] (N=82).
OOD comparison methods. The RF-Deep classifier was compared against standard OOD methods including MaxSoftmax [hendrycks17baseline], MaxLogits [hendrycks2022logits], energy [liu2020energy], and entropy measures. In addition, a RF-radiomics classifier was created using 293 IBSI-compliant radiomic features [zwanenburg2020image] extracted from detected tumor regions using PyCERR radiomics library [deasy2003cerr, apte2018extension]. In order to prevent accuracy degradation from correlated features, recursive feature elimination was employed. RF used 1000 trees, max depth of 20, and balanced weights.
OOD experiment protocol. RF-based approaches used a fixed 40/60 patient-level train–test split, repeated over 100 seeds. Other baselines used the full cohorts as they require no auxiliary data.
Evaluation metrics. We report AUROC to measure ID–OOD separability [davis2006relationship] and FPR95 as a threshold-based metric [liang2017enhancing].
@@ 3 Results
Table 1: Out-of-distribution (OOD) detection performance comparing RF-Deep and other methods. Results are reported as AUROC ($\uparrow$) and FPR95 ($\downarrow$), wiith best values per dataset highlighted in bold.
Method Pulmonary Embolism COVID-19 Kidney Cancer Abdomen
AUROC ($\uparrow$) FPR95 ($\downarrow$) AUROC ($\uparrow$) FPR95 ($\downarrow$) AUROC ($\uparrow$) FPR95 ($\downarrow$) AUROC ($\uparrow$) FPR95 ($\downarrow$)
@@
Our results demonstrate that RF-Deep provides a robust scan-level OOD detection for lung cancer segmentation, outperforming established approaches including MaxSoftmax and MaxLogits ( 1). SimMIM-pretrained encoder features, paired with RF, consistently yielded higher AUROC and lower FPR95 on our lung cancer segmentation task, with large margin gains in the abdominal cohorts (FPR95 $\leq$ 0.1, second best is 6 points away). Fig. 2 visually shows RF-Deep’s ability to detect scans from completely different anatomical sites and different disease cases, with some limitations. In pulmonary embolism cases, we observed that the segmentation model often highlights tumor-like structures rather than the emboli; nevertheless, our contextual features enable RF-Deep to correctly classify these scans as OOD most of the times, surpassing non-contextual approaches.
Additionally, we performed t-SNE clustering of the features [masarczyk2024tunnel] across all cohorts (Fig. 3(a)) and observed partially distinct clusters for in-distribution (ID) and OOD cohorts, making them favourably separable with the random forest classifier. Fig. 3(b) and Fig. 3(c) show the SHAP analysis on the RF-Deep providing partial interpretability in identifying features that are most influential in classifying a scan as OOD. Finally, stage-wise ablation (Fig. 3(d)), wherein only features corresponding to a particular stage of the transformer were used for training RF-Deep, demonstrated that mid- and early-deep level Swin Transformer features are most effective, aligning with our SHAP findings.